OpenStudy (anonymous):

Find S15 for the series -1 + -3 + -5 + -7 +..

OpenStudy (anonymous):

t1= -1 d= -2 n=15 t15 = -1 + (15-1)-2 t15 = -1 + (14)-2 t15 = -1 + -28 t15 = -29 S15 = 15 (-1 + -29) / 2 = -225 is this right?